suppose you are starting an office cleaning service. You have spent $315 on equipment. To clean an office, you use $4 worth of supplies you. You charge 25$ per office. How many offices must you clean to break even
Answer by stanbon(57412)

suppose you are starting an office cleaning service. You have spent $315 on equipment. To clean an office, you use $4 worth of supplies you. You charge 25$ per office. How many offices must you clean to break even
---
E(x) = 315+4x where x is the number of offices you clean and E is expense.
---
R(x) = 25x
------
Equation:
Solve E(x) = R(x)
315+4x = 25x
21x = 315
x = 15 offices
==========
Cheers,
Stan H.